OpenStudy (anonymous):

A BALL IS THROWN UPWARDS ON THE MOON WITH AN INITIAL SPEED OF 35m/s .TAKE g ON MOON =1.6 CALCULATE (a)the time taken to reach the maximum height (b)the maximum height reached by the ball (c) its velocity 3 seconds after it has been thrown

OpenStudy (anonymous):

(a) v(t)=at+v(initial)=0 35+(-1.6)t=0 then you get t

OpenStudy (anonymous):

for (b) v(3)=a*3+v(initial)=-1.6*3+35